# Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in

### v2

@@ 2 The singlet-doublet dark matter model
In this case, the SM is extended with a fermion doublet $\Psi=(\psi^{0}\leavevmode\nobreak\ \leavevmode\nobreak\ \psi^{-})^{T}$ and a singlet fermion $\chi$, which are stabilized by a discrete symmetry $\mathcal{Z}_{2}$, under which they are odd, and all other SM particles are charged even [6].
@@
$M_{\chi_{0}}$ $=$ $M_{\Psi}\cos^{2}\theta+\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\sin^{2}\theta,\leavevmode\nobreak\ \leavevmode\nobreak\$
@@ 3 Dark matter phenomenology
In this section, we systematically explore the relic density parameter space for SDDM, considering annihilation, co-annihilation, co-scattering, and freeze-in, depending on the $\sin\theta$ and $\Delta{M}$. We divide the parameter space into four regimes: (i) thermal DM relic considering annihilation and co-annihilation only, (ii) thermal DM relic via conversion-driven processes processes, (iii) dark matter production via freeze-in mechanism, and (iv) dark matter production via SuperWIMP mechanism.
@@ 3.1 Thermal production of dark matter
$\Omega_{\chi_{1}}h^{2}\equiv\Omega_{\rm DM}h^{2}=\frac{1.09\times 10^{9}{\rm GeV}^{-1}}{\sqrt{g_{*}}M_{\rm Pl}}\left[\int_{x_{f}}^{\infty}dx\leavevmode\nobreak\ \frac{\langle\sigma v\rangle_{\rm eff}}{x^{2}}\right]^{-1}.$ (3.4)
@@ 3.1.1 Thermal DM relic in the large mixing regime via annihilation and co-annihilation
$\rho_{1}=\frac{\Omega_{1s}h^{2}}{\Omega_{2s}h^{2}(\rm no\leavevmode\nobreak\ coscattering)},$ (3.9)
where $\Omega_{1s}h^{2}$ is the relic density obtained by solving Eq. (3.1), and $\Omega_{2s}h^{2}$(no co-scattering) is calculated by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) excluding co-scattering processes [^4] We use micrOMEGAs 6.1.15[53, 42] to solve the Boltzmann equations, where darkOmega function is used to estimate $\Omega_{1s}h^{2}$ and darkOmegaN is used to estimate $\Omega_{2s}h^{2}$(no co-scattering) while excluding the co-scattering processes. This can be achieved by using the flag ExcludedForNDM="2010", while using darkOmegaN function.. In particular we have switched off $\langle\sigma_{2010}v\rangle n^{\rm eq}_{\rm SM}$ in Eq. (3.8). In Eq. (3.9), $\rho_{1}=1$ implies that all the dark sector particles decouple at the same epoch. On the other hand, $\rho_{1}<1$ implies that the dark sector particles are decoupling at different epochs. When $\sin\theta$ is small, the singlet decouples earlier than the doublet and produces a relatively larger relic in absence of the co-scattering processes. This implies that $\Omega_{2s}h^{2}({\rm no\leavevmode\nobreak\ coscatttering})>\Omega_{1s}h^{2}$ Therefore, one expects $\rho_{1}\leq 1$.
@@ 3.1.2 Thermal dark matter relic via the conversion-driven processes in the small $\sin\theta$ regime
As discussed in section 3.1.1, in the large SD mixing angle limit ($\mathcal{O}(10^{-4})<\sin\theta<0.5$) only annihilation and co-annihilation are the dominating processes that decide the DM relic. Had we been added the co-scattering processes to Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), the result would not have changed. This can be understood by analyzing Eq. (3.8), where the second term contributes to the effective co-scattering processes. At this point we note that in the large $\sin\theta$ limit, the impact of the co-scattering processes are similar in magnitude to that of the decay and inverse decay processes given in Eq. (3.8). The large co-scattering processes can interchange the sector 1 particles to sector 2 particles efficiently. If the sector 2 particles are already in equilibrium, then they can annihilate to SM particles and deplete the dark matter number density through annihilation and co-annihilation processes $\psi^{+}\psi^{-}\rightarrow\rm\rm SM\leavevmode\nobreak\ SM$ which is independent of the SD mixing angle. On the other hand, if the singlet and doublet decoupled already (the processes 1100,1200 and 2200 decoupled) then the large conversion-driven processes can not alter the total relic, $Y_{1}+Y_{2}$, even if these processes remain in equilibrium for a longer period. As a result, the impact of the co-scattering processes is not felt in the large mixing angle limit. In this case, the decoupling of particles in both sectors happens in the same epoch and the Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) simplify to the one-sector BE as given in Eq. (3.1).
@@
BP1 ($\sin\theta=0.119$): In the left panel of Fig. 2, we depict the evolution of the abundances of both dark sector species as a function of temperature for mixing angles: $\sin\theta=1.19\times 10^{-1}$. The dark matter mass is fixed at $146.48\leavevmode\nobreak\ \mathrm{GeV}$, and $\Delta{M}=10.5\leavevmode\nobreak\ \mathrm{GeV}$, selected to obtain the correct relic abundance. The comparison among relevant interaction rates is displayed in the right panel of Fig. 2. Here, we observe that the annihilation rate of the singlet component (1100 processes) falls below the Hubble rate first, followed by the decoupling of co-annihilation processes (1200 processes). Both sectors remain in chemical equilibrium down to $T\sim 6\leavevmode\nobreak\ \mathrm{GeV}$, as seen in the left panel of Fig. 2. On the other hand, the rate of conversion-driven processes (decay, inverse decay and co-scattering) remain significantly larger than the Hubble rate during this period. These processes can efficiently deplete the number density of dark matter by efficiently converting the sector 1 particles to sector 2 particles followed by annihilation to SM particles (2200 processes). Once the 2200 processes decouple (shown by the red line in right panel of Fig. 2) around $T\sim 5$ GeV, the number density is settled to its final value as seen in the left panel of Fig. 2.
@@
BP2 ($\sin\theta=1.19\times 10^{-5}$): In the left panel of Fig. 3, we depict the evolution of the abundances of both dark sector species as a function of temperature for mixing angles: $\sin\theta=1.19\times 10^{-5}$. The DM mass is taken to be 146.48 GeV, and $\Delta{M}=8.84\leavevmode\nobreak\ \mathrm{GeV}$ to get the correct relic density. Since the rate of sector 1 annihilation processes (1100) is proportional to $\sin^{4}\theta$, these processes are negligible in comparison to co-annihilation processes (1200) for this BP as the rate of co-annihilation processes are proportional to $\sin^{2}\theta$. As shown in the right panel of Fig. 3, the co-annihilation processes decouple at approximately $T\sim 20\leavevmode\nobreak\ \mathrm{GeV}$. As a result the chemical decoupling of sector 1 particles happens with a relatively larger abundance, while sector 2 particles remain in equilibrium due to their gauge interactions, as shown in the left panel of Fig. 3. Note that in the small mixing angle limit, the decay and inverse decay processes are subdominant in comparison to co-scattering processes. This is evident from the right panel of Fig. 3. It is worth mentioning that, in this case, processes involving light SM fermions contribute to the co-scattering processes significantly, e.g. $\chi_{1}e^{-}\rightarrow\psi^{-}\nu_{e},\chi_{1}\nu_{e}\rightarrow\psi^{+}e^{-}$ etc. In the presence of efficient co-scattering, sector 1 particle can upscatter with SM particles into sector 2, followed by annihilation to SM via 2200 processes. Once the 2200 processes decouple, the relic settles to its final value even if the rate of co-scattering processes continue to be in equilibrium for a longer period. For this BP the 2200 processes decouple at $T\sim 5$ GeV (see red line in the right panel of Fig. 3.). As a result the relic settles around $T\sim 5$ GeV as seen from the left panel of Fig. 3.
@@
Figure 4: Variation of DM relic density w.r.t. DM mass with the fixed mass splitting of $1\leavevmode\nobreak\ \rm GeV\leq\Delta{M}\leq 5$ GeV for $10^{-2}\leq\sin\theta\leq 10^{-1}$ [left] and $10^{-5}\leq\sin\theta\leq 10^{-4}$ [right]. For the blue colored points, we have solved Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), by switching off the co-scattering processes, while for the red points we have considered all the processes given in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2).
@@
Figure 5: Variation of DM relic density w.r.t DM mass with the fixed mass splitting of $5\leavevmode\nobreak\ {\rm GeV}\leq\Delta{M}\leq 10$ GeV for $10^{-2}\leq\sin\theta\leq 10^{-1}$ [left] and $10^{-5}\leq\sin\theta\leq 10^{-4}$ [right]. For the blue colored points, we have solved Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), by switching off the co-scattering processes, while for the red points we have considered all the processes given in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2).
@@
$\rho_{2}=\frac{\Omega_{2s}h^{2}}{\Omega_{2s}h^{2}({\rm no\leavevmode\nobreak\ coscattering})},$ (3.12)
@@ 3.2 Dark matter production via freeze-in mechanism
In Sections 3.1.1, and 3.1.2, we saw that the singlet is brought to equilibrium for the $\sin\theta\gtrsim\mathcal{O}(10^{-7})$ depending on the DM mass. This implies that for $\sin\theta\gtrsim\mathcal{O}(10^{-7})$, the relic density of SDDM will be obtained via the freeze-out mechanism. If the SD mixing is further reduced, the DM relic density can no longer be determined through the conventional freeze-out mechanisms. However, it can be produced through the freeze-in mechanism. In this regime, we take into account all the $2\rightarrow 2$ processes along with decay and inverse decay processes and solve Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2) with initial conditions, $Y_{2}=Y_{2}^{\rm eq}$ and $Y_{1}=0$. To estimate the final DM relic density, we used $\Omega_{\rm DM}h^{2}=0.12\left(\frac{(Y_{1}+Y_{2})_{T\rightarrow 0}}{4.26\times 10^{-10}}\right)\left(\frac{M_{\rm DM}}{{1\leavevmode\nobreak\ \rm GeV}}\right)$.
@@ 4 Direct detection of dark matter
Different coupling strengths between DM and light quarks are given by [55, 56] as $f^{p}_{Tu}=0.020\pm 0.004,\leavevmode\nobreak\ f^{p}_{Td}=0.026\pm 0.005,\leavevmode\nobreak\ f^{p}_{Ts}=0.014\pm 0.062,\leavevmode\nobreak\ f^{n}_{Tu}=0.020\pm 0.004,\leavevmode\nobreak\ f^{n}_{Td}=0.036\pm 0.005,\leavevmode\nobreak\ f^{n}_{Ts}=0.118\pm 0.062$. The coupling of DM with the gluons in target nuclei are parameterized by [57] $f^{p,n}_{TG}=1-\sum_{q=u,d,s}f^{p,n}_{Tq}$. Similarly, the SI DM-nucleon scattering cross-section through Z mediation is given by [58, 59]:
@@ 6 Conclusions
7 Acknowledgments
P.K.P. would like to acknowledge the Ministry of Education, Government of India, for providing financial support for his research via the Prime Minister’s Research Fellowship (PMRF) scheme. The work of N.S. is supported by the Department of Atomic Energy-Board of Research in Nuclear Sciences, Government of India (Ref. Number: 58/14/15/2021- BRNS/37220). We would like to thank Alexander Pukhov for his assistance on $\tt micrOMEGAs$.
@@ Appendix C Decay rates
${\rm where,\leavevmode\nobreak\ }\Lambda(x,y,z)$ $=$ $x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$ (C.2)

### v3

@@ 2 The singlet-doublet dark matter model
In this case, the SM is extended with a fermion doublet $\Psi=(\psi^{0}\penalty 10000\ \penalty 10000\ \psi^{-})^{T}$ and a singlet fermion $\chi$, which are stabilized by a discrete symmetry $\mathcal{Z}_{2}$, under which they are odd, and all other SM particles are charged even [6].
@@
$M_{\chi_{0}}$ $=$ $M_{\Psi}\cos^{2}\theta+\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\sin^{2}\theta,\penalty 10000\ \penalty 10000\$
@@ 3 Dark matter phenomenology
In this section, we systematically explore the relic density parameter space for SDDM, considering annihilation, co-annihilation, co-scattering, and freeze-in, depending on the $\sin\theta$ and $\Delta{M}$. We divide the parameter space into four regimes: (i) thermal DM relic considering annihilation and co-annihilation only, (ii) thermal DM relic via conversion-driven processes, (iii) dark matter production via freeze-in mechanism, and (iv) dark matter production via SuperWIMP mechanism.
@@ 3.1 Thermal production of dark matter
$\Omega_{\chi_{1}}h^{2}\equiv\Omega_{\rm DM}h^{2}=\frac{1.09\times 10^{9}{\rm GeV}^{-1}}{\sqrt{g_{*}}M_{\rm Pl}}\left[\int_{x_{f}}^{\infty}dx\penalty 10000\ \frac{\langle\sigma v\rangle_{\rm eff}}{x^{2}}\right]^{-1}.$ (3.4)
@@ 3.1.1 Thermal DM relic in the large mixing regime via annihilation and co-annihilation
$\rho_{1}=\frac{\Omega_{1s}h^{2}}{\Omega_{2s}h^{2}(\rm no\penalty 10000\ coscattering)},$ (3.9)
where $\Omega_{1s}h^{2}$ is the relic density obtained by solving Eq. (3.1), and $\Omega_{2s}h^{2}$(no co-scattering) is calculated by solving Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) excluding co-scattering processes [^4] We use micrOMEGAs 6.1.15[53, 42] to solve the Boltzmann equations, where darkOmega function is used to estimate $\Omega_{1s}h^{2}$ and darkOmegaN is used to estimate $\Omega_{2s}h^{2}$(no co-scattering) while excluding the co-scattering processes. This can be achieved by using the flag ExcludedForNDM="2010", while using darkOmegaN function.. In particular we have switched off $\langle\sigma_{2010}v\rangle n^{\rm eq}_{\rm SM}$ in Eq. (3.8). In Eq. (3.9), $\rho_{1}=1$ implies that all the dark sector particles decouple at the same epoch. On the other hand, $\rho_{1}<1$ implies that the dark sector particles are decoupling at different epochs. When $\sin\theta$ is small, the singlet decouples earlier than the doublet and produces a relatively larger relic in absence of the co-scattering processes. This implies that $\Omega_{2s}h^{2}({\rm no\penalty 10000\ coscatttering})>\Omega_{1s}h^{2}$ Therefore, one expects $\rho_{1}\leq 1$.
@@ 3.1.2 Thermal dark matter relic via the conversion-driven processes in the small $\sin\theta$ regime
As discussed in section 3.1.1, in the large SD mixing angle limit ($\mathcal{O}(10^{-4})<\sin\theta<0.5$) only annihilation and co-annihilation are the dominating processes that decide the DM relic. Had we been added the co-scattering processes to Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), the result would not have changed. This can be understood by analyzing Eq. (3.8), where the second term contributes to the effective co-scattering processes. At this point we note that in the large $\sin\theta$ limit, the impact of the co-scattering processes are similar in magnitude to that of the decay and inverse decay processes given in Eq. (3.8). The large co-scattering processes can interchange the sector 1 particles to sector 2 particles efficiently. If the sector 2 particles are already in equilibrium, then they can annihilate to SM particles and deplete the dark matter number density through annihilation and co-annihilation processes $\psi^{+}\psi^{-}\rightarrow\rm\rm SM\penalty 10000\ SM$ which is independent of the SD mixing angle. On the other hand, if the singlet and doublet decoupled already (the processes 1100,1200 and 2200 decoupled) then the large conversion-driven processes can not alter the total relic, $Y_{1}+Y_{2}$, even if these processes remain in equilibrium for a longer period. As a result, the impact of the co-scattering processes is not felt in the large mixing angle limit. In this case, the decoupling of particles in both sectors happens in the same epoch and the Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) simplify to the one-sector BE as given in Eq. (3.1).
@@
BP1 ($\sin\theta=0.119$): In the left panel of Fig. 2, we depict the evolution of the abundances of both dark sector species as a function of temperature for mixing angles: $\sin\theta=1.19\times 10^{-1}$. The dark matter mass is fixed at $146.48\penalty 10000\ \mathrm{GeV}$, and $\Delta{M}=10.5\penalty 10000\ \mathrm{GeV}$, selected to obtain the correct relic abundance. The comparison among relevant interaction rates is displayed in the right panel of Fig. 2. Here, we observe that the annihilation rate of the singlet component (1100 processes) falls below the Hubble rate first, followed by the decoupling of co-annihilation processes (1200 processes). Both sectors remain in chemical equilibrium down to $T\sim 6\penalty 10000\ \mathrm{GeV}$, as seen in the left panel of Fig. 2. On the other hand, the rate of conversion-driven processes (decay, inverse decay and co-scattering) remain significantly larger than the Hubble rate during this period. These processes can efficiently deplete the number density of dark matter by efficiently converting the sector 1 particles to sector 2 particles followed by annihilation to SM particles (2200 processes). Once the 2200 processes decouple (shown by the red line in right panel of Fig. 2) around $T\sim 5$ GeV, the number density is settled to its final value as seen in the left panel of Fig. 2.
@@
BP2 ($\sin\theta=1.19\times 10^{-5}$): In the left panel of Fig. 3, we depict the evolution of the abundances of both dark sector species as a function of temperature for mixing angles: $\sin\theta=1.19\times 10^{-5}$. The DM mass is taken to be 146.48 GeV, and $\Delta{M}=8.84\penalty 10000\ \mathrm{GeV}$ to get the correct relic density. Since the rate of sector 1 annihilation processes (1100) is proportional to $\sin^{4}\theta$, these processes are negligible in comparison to co-annihilation processes (1200) for this BP as the rate of co-annihilation processes are proportional to $\sin^{2}\theta$. As shown in the right panel of Fig. 3, the co-annihilation processes decouple at approximately $T\sim 20\penalty 10000\ \mathrm{GeV}$. As a result the chemical decoupling of sector 1 particles happens with a relatively larger abundance, while sector 2 particles remain in equilibrium due to their gauge interactions, as shown in the left panel of Fig. 3. Note that in the small mixing angle limit, the decay and inverse decay processes are subdominant in comparison to co-scattering processes. This is evident from the right panel of Fig. 3. It is worth mentioning that, in this case, processes involving light SM fermions contribute to the co-scattering processes significantly, e.g. $\chi_{1}e^{-}\rightarrow\psi^{-}\nu_{e},\chi_{1}\nu_{e}\rightarrow\psi^{+}e^{-}$ etc. In the presence of efficient co-scattering, sector 1 particle can upscatter with SM particles into sector 2, followed by annihilation to SM via 2200 processes. Once the 2200 processes decouple, the relic settles to its final value even if the rate of co-scattering processes continue to be in equilibrium for a longer period. For this BP the 2200 processes decouple at $T\sim 5$ GeV (see red line in the right panel of Fig. 3.). As a result the relic settles around $T\sim 5$ GeV as seen from the left panel of Fig. 3.
@@
Figure 4: Variation of DM relic density w.r.t. DM mass with the fixed mass splitting of $1\penalty 10000\ \rm GeV\leq\Delta{M}\leq 5$ GeV for $10^{-2}\leq\sin\theta\leq 10^{-1}$ [left] and $10^{-5}\leq\sin\theta\leq 10^{-4}$ [right]. For the blue colored points, we have solved Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), by switching off the co-scattering processes, while for the red points we have considered all the processes given in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2).
@@
Figure 5: Variation of DM relic density w.r.t DM mass with the fixed mass splitting of $5\penalty 10000\ {\rm GeV}\leq\Delta{M}\leq 10$ GeV for $10^{-2}\leq\sin\theta\leq 10^{-1}$ [left] and $10^{-5}\leq\sin\theta\leq 10^{-4}$ [right]. For the blue colored points, we have solved Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2), by switching off the co-scattering processes, while for the red points we have considered all the processes given in Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2).
@@
$\rho_{2}=\frac{\Omega_{2s}h^{2}}{\Omega_{2s}h^{2}({\rm no\penalty 10000\ coscattering})},$ (3.12)
@@ 3.2 Dark matter production via freeze-in mechanism
In Sections 3.1.1, and 3.1.2, we saw that the singlet is brought to equilibrium for the $\sin\theta\gtrsim\mathcal{O}(10^{-7})$ depending on the DM mass. This implies that for $\sin\theta\gtrsim\mathcal{O}(10^{-7})$, the relic density of SDDM will be obtained via the freeze-out mechanism. If the SD mixing is further reduced, the DM relic density can no longer be determined through the conventional freeze-out mechanisms. However, it can be produced through the freeze-in mechanism. In this regime, we take into account all the $2\rightarrow 2$ processes along with decay and inverse decay processes and solve Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2) with initial conditions, $Y_{2}=Y_{2}^{\rm eq}$ and $Y_{1}=0$. To estimate the final DM relic density, we used $\Omega_{\rm DM}h^{2}=0.12\left(\frac{(Y_{1}+Y_{2})_{T\rightarrow 0}}{4.26\times 10^{-10}}\right)\left(\frac{M_{\rm DM}}{{1\penalty 10000\ \rm GeV}}\right)$.
@@ 4 Direct detection of dark matter
Different coupling strengths between DM and light quarks are given by [55, 56] as $f^{p}_{Tu}=0.020\pm 0.004,\penalty 10000\ f^{p}_{Td}=0.026\pm 0.005,\penalty 10000\ f^{p}_{Ts}=0.014\pm 0.062,\penalty 10000\ f^{n}_{Tu}=0.020\pm 0.004,\penalty 10000\ f^{n}_{Td}=0.036\pm 0.005,\penalty 10000\ f^{n}_{Ts}=0.118\pm 0.062$. The coupling of DM with the gluons in target nuclei are parameterized by [57] $f^{p,n}_{TG}=1-\sum_{q=u,d,s}f^{p,n}_{Tq}$. Similarly, the SI DM-nucleon scattering cross-section through Z mediation is given by [58, 59]:
@@ Appendix C Decay rates
${\rm where,\penalty 10000\ }\Lambda(x,y,z)$ $=$ $x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$ (C.2)
@@
Acknowledgments
P.K.P. would like to acknowledge the Ministry of Education, Government of India, for providing financial support for his research via the Prime Minister’s Research Fellowship (PMRF) scheme. The work of N.S. is supported by the Department of Atomic Energy-Board of Research in Nuclear Sciences, Government of India (Ref. Number: 58/14/15/2021- BRNS/37220). We would like to thank Alexander Pukhov for his assistance on $\tt micrOMEGAs$. We would like acknowledge the anonymous referee of JCAP for his/her constructive comments which improved the manuscript.